### v3

# The GUS Framework: Benchmarking Social Bias Classification with Discriminative (Encoder-Only) and Generative (Decoder-Only) Language Models

Abstract
The detection of social bias in text is a critical challenge, particularly due to the limitations of binary classification methods. These methods often oversimplify nuanced biases, leading to high emotional impact when content is misclassified as either “biased” or “fair.” To address these shortcomings, we propose a more nuanced framework that focuses on three key linguistic components underlying social bias: Generalizations, Unfairness, and Stereotypes (the GUS framework). The GUS framework employs a semi-automated approach to create a comprehensive synthetic dataset, which is then verified by humans to maintain ethical standards. This dataset enables robust multi-label token classification. Our methodology, which combines discriminative (encoder-only) models and generative (auto-regressive large language models), identifies biased entities in text. Through extensive experiments, we demonstrate that encoder-only models are effective for this complex task, often outperforming state-of-the-art methods, both in terms of macro and entity-wise F1-score and Hamming loss. These findings can guide the choice of model for different use cases, highlighting the GUS framework’s effectiveness in capturing explicit and implicit biases across diverse contexts, and offering a pathway for future research and applications in various fields. GUS resources can be found here: https://huggingface.co/collections/ethical-spectacle/gus-net-social-bias-ner-66edfe93801ea45d7a26a10f.
[…]
1 Introduction
The importance of social bias analysis in natural language processing (NLP) is increasing [1], particularly as communication increasingly relies on Large Language Models (LLMs) [2] across various domains such as education [3] and business [4]. Social bias can influence public perception and decision-making, often subtly reinforcing stereotypes or discriminatory practices. While explicit bias, which refers to overt prejudice or favoritism, is relatively easy to identify, implicit bias involves more subtle and often unconscious associations or attitudes[5]. Detecting and mitigating implicit bias in the text is significantly more challenging, as perceptions of bias can vary greatly depending on the context, including the perspectives of readers and speakers.
For example, consider the phrase, ”Hard-working immigrants contribute significantly to society.” To some, this statement may appear positive, acknowledging the effort and diligence of immigrants. However, from another perspective, it might be perceived as implicitly biased, suggesting that immigrants are expected to work harder than others to be valued or accepted. This subtle implication can reinforce stereotypes that separate immigrants from native citizens, placing an undue burden of proof on their worthiness. Such subjectivity highlights the complexity of implicit bias detection, making it a critical area of research within NLP [6, 7, 8].
In state-of-the-art research, much of the focus remains on detecting biases at the sentence level [9, 10, 11]. While this approach is useful, there is a growing need for more granular bias detection. Token classification [12], or named-entity recognition (NER), enables the identification of specific words or tokens contributing to biased sentiment. This approach is often more suitable than sequence classification, which involves a higher level of abstraction that can lead to ambiguity and disagreement. However, research on linguistic bias detection at the granular level remains limited. For instance, Nbias [8] focuses on a single entity like ”BIAS,” but broader and more comprehensive frameworks are still lacking. This can be attributed to the resource-intensive nature of annotating and reviewing input text with token classes. For example, finding agreements between annotators is practical when input text is assigned a single label, as in binary classification. However, enforcing consistency of entity usage and boundaries is exponentially more difficult in human token-labeling due to personal tendencies and a less controlled environment than can now be created with LLM agents for labeling.
Modern named-entity recognition methods vary, often specifically designed for the environment they’ll be deployed. For example, social media posts tend to be unstructured or grammatically nuanced, incorporating pragmatics that diverge from conventional natural language. To handle this, recent methods incorporate data augmentation, semantic transformation, and multiple models/architectures such as BERT and LSTM to build a robust framework [13, 14, 15]. However, named-entity recognition training methods have also shown susceptibility to bias in the training data, often underperforming on specific demographics across datasets. Further, de-biasing of training data isn’t likely to improve the performance on under-represented demographics, instead fair corpus distribution is imperative [16].
One of the primary challenges in this domain is the scarcity of suitable datasets for training token classification models. While implicit bias can manifest in subtle forms, such as word choice, narrative framing, or the omission of certain viewpoints, capturing these patterns in a dataset presents unique challenges. Existing datasets, such as the Media Bias Annotation Dataset (MBIC) [17] and Bias Annotations By Experts (BABE) [18], rely on human annotators. However, the annotation process varies in sophistication. For example, MBIC outsources annotations to non-experts via Amazon Mechanical Turk, while BABE employs in-house expert annotators and ensures consensus among them. Although useful, both approaches come with trade-offs in terms of cost, dataset size, or label quality, which are further exacerbated as task complexity and granularity increase. As a result, it has become conventional to automate the extrapolation of human-labeled data, incorporating humans-in-the-loop for validation.
In response to these challenges, this paper introduces the Generalizations, Unfairness, and Stereotypes Dataset (GUS Dataset), which captures three types of bias commonly discussed in legal and psychological literature [19, 20, 21, 22]: Generalizations, Unfairness, and Stereotypes. The GUS Dataset leverages generative AI and automated agents to construct an optimal dataset for bias detection. This approach improves upon traditional methods by combining LLM reasoning with human-in-the-loop supervision, resulting in more accurate, granular, and comprehensive social bias patterns and labels across domains.
The main contributions of this paper are as follows:
1.
We generate a corpus of 3,739 text snippets depicting biases across varied domains such as religious bias, racial bias, political bias, and others. The snippets are labeled through an AI-human collaborative process, where a team of LLM agents first identifies biased words up to 4/5-grams, followed by a review by five human annotators to verify the accuracy of the labels.
2.
We format the annotations into a list of lists where each input token can be assigned multiple classes, for a multi-label token classification problem. To the best of our knowledge, this is the first work to implement data and methods for multi-label token classification with entities such as Generalizations (B-GEN, I-GEN), Unfairness (B-UNFAIR, I-UNFAIR), Stereotypes (B-STEREO, I-STEREO), or Neutral (O).
3.
We benchmark two families of models: encoder-only models (e.g., BERT-like architectures) and decoder-only auto-regressive models (e.g., LLMs), for multi-label named-entity recognition, trained on the GUS Dataset. This is the first work to provide token-level multi-label bias detection.
4.
We conduct experiments to demonstrate the contributions of our methods in relation to existing approaches, showcasing improvements in accuracy, F1-score, and the depth of bias detection.
Empirical results on two model families show that encoder-only models are more effective at identifying nuanced biases, often outperforming generative models in terms of accuracy, F1-score, and Hamming loss, while maintaining computational efficiency. However, further investigation is needed to assess the reasoning capabilities of these models, as this aspect was not explicitly evaluated in our study.
2 Related Works
The detection of social bias in natural language processing (NLP) is a critical area of research, particularly given the increasing use of large language models (LLMs) across various domains [23, 24, 2, 3, 25, 4, 26]. Traditional techniques for bias detection often rely on human annotators to label datasets. While this approach has been instrumental in creating foundational resources, it is inherently limited by the annotators’ resources and expertise. This limitation often leads to datasets/frameworks that contain a narrow understanding of bias, especially in regards to implicit biases that are subtle and context-dependent [1, 6, 7].
2.1 Ethical Dataset Construction
The construction of ethical datasets for bias detection is essential for ensuring comprehensive and fair analyses. Existing datasets often suffer from limitations in scope, failing to encompass the broad spectrum of biases and perspectives necessary for effective bias detection. For example, the Dbias model [27] utilized the MBIC dataset, which consists of a relatively small number of sentences, restricting the model’s ability to generalize across different domains and types of bias. Although the Nbias framework [8] expanded the use of named-entity recognition (NER) by introducing the entity “BIAS” it still primarily addressed explicit biases and overlooked the structural elements of implicit bias, such as stereotypes and generalizations.
Moreover, studies that emphasize robust annotations often rely on human judgment, which can lead to a lack of diversity in viewpoints necessary to capture the nuance of implicit bias [18]. This reliance on human annotators may also perpetuate the biases present in society, resulting in datasets that do not adequately represent the full range of perspectives. Thus, there is a pressing need for more diverse and comprehensive datasets that can capture implicit biases in language.
2.2 Bias Detection
Traditional methods typically focus on explicit bias, which is easier to define and identify, while neglecting the subtler forms of bias that may influence public perception and decision-making. Implicit bias can manifest through word choice, framing, and the omission of certain viewpoints, making it challenging to detect using conventional approaches [1, 6].
Existing frameworks, such as Dbias and Nbias, have made strides in bias detection but still focus primarily on explicit biases, leaving a gap in the understanding of how implicit biases operate [27, 8]. Additionally, the datasets used for these frameworks often lack the necessary diversity of perspectives, limiting their effectiveness in identifying implicit biases. In contrast, our proposed approach leverages generative AI and automated agents to construct a more comprehensive dataset. By utilizing synthetic data generated by these agents, we enhance the training of the pre-trained model BERT for multi-label token classification. This innovative methodology not only improves the specificity and depth of bias detection but also addresses the limitations of existing datasets, paving the way for more accurate and nuanced understanding of biases in various texts.
3 GUS-Net Framework for Social Bias Detection
Problem Definition We aim to detect and classify words and phrases in text that express social bias—specifically generalizations, unfair language, and stereotypes—using a named entity recognition approach with multi-label classification. This is a token-level multi-label classification task, where each sentence is annotated with one or more labels to facilitate the identification of biased entities across token sequences.
3.1 Data Preparation
Team Formation
The annotation team consisted of five members, each with a background in computational linguistics or social sciences. All team members held at least a bachelor’s degree in science, ensuring a strong foundation in both technical and domain-specific knowledge. Guidelines were provided to ensure consistency in the annotation process, including detailed instructions on entity definitions, labeling conventions, and conflict resolution protocols. Regular meetings were held to address ambiguities and maintain alignment across the team.
Synthetic Data
To construct a comprehensive dataset for social bias classification, we leverage modern synthetic training data techniques to avoid the labor-intensive and potentially subjective human annotation process [28, 29, 30]. Our approach consists of two main steps: synthetic annotation and corpus generation using a language model for enhanced coverage and structural consistency [31, 32, 33]. Figure 2 illustrates the dataset generation pipeline, and Table 1 shows the arguments combined to form prompts across domains.
Human-in-the-Loop Review and Pre-processing (Corpus)
Before annotation, the synthetic corpus underwent qualitative review by human evaluators. Reviewers examined 5-10 samples per bias type and demographic group, ensuring that generated statements were realistic and free of counterfactual stereotypes. Reviewers were also allowed to remove unrealistic cases or refine prompt categories where necessary. However, such instances were rare, confirming that Mistral-7B effectively generated relevant bias-based statements. Finally, a quantitative distribution analysis of the prompt arguments ensured balanced representation across bias types and demographic groups. Once approved, the dataset proceeded to the annotation stage.
| Type Of Bias | Target | Statement Type | Sentiment |
| Racial | [white people, black people, asian people, hispanic people, indigenous people, …] | ||
| Religious | [christians, muslims, jewish people, hindus, buddhists, sikhs, atheists, agnostics, …] | ||
| Gender | [men, women, boys, girls, females, males, non-binary people, …] | ||
| Age | [children, teenagers, young people, middle aged people, old people, …] | ||
| Nationality | [immigrants, refugees, people from developing countries, people from Western countries, …] | ||
| Sexuality | [straight people, gay people, bisexual people, asexual people, LGBTQIA+ people, …] | ||
| Socioeconomic | [working class people, middle class people, upper class people, poor people, rich people, …] | ||
| Educational | [uneducated people, highly educated people, people with non-traditional education, …] | ||
| Disability | [people with physical disabilities, people in wheelchairs, people with mental disabilities, …] | ||
| Political | [republicans, democrats, independents, conservatives, liberals, progressives, …] | ||
| Physical | [tall people, short people, fat people, skinny people, ugly people, hot people, …] | Stereotypes, Unfair Generalizations, False Assumptions, Discriminatory Language, Offensive Implications | Positive, Negative |
Data Annotation
The data annotation process involves several systematic steps to ensure accuracy and consistency in labeling the generated sentences. We utilized GPT-4o and the Stanford DSPy framework [34] to annotate the generated sentences with entity labels, following methodologies in recent literature for annotation with agents [35]. As shown in Figure 2, the annotation can be summarized in the following steps:
Preparation of annotation: We recompiled a DSPy agent for each entity type (e.g., generalization, unfairness, stereotype) as an agent to streamline the annotation process. Each agent contains the entity definition and four examples of correct annotations. The prompts sent to an agent include the definition of the target entity along with the curated examples. Using few-shot prompting helps the agent understand the context and provides guidance for accurate labeling [36, 37, 38, 39]. Their definitions are provided in Table 2.
| Class | Prompt Definition | Entity Label |
| Generalizations | Any broad generalization of a group or ubiquitous classifiers, including adjectives and descriptors. | B-GEN, I-GEN |
| Unfairness | Any harsh or unjust characterization or offensive language. | B-UNFAIR, I-UNFAIR |
| Stereotypes | Any statement (multiple words) that contains a stereotype targeting a group of people, both explicitly and unconsciously. | B-STEREO, I-STEREO |
| Neutral | - | O |
Annotation by agents: Each generated sentence is processed by an agent supported by an LLM for a single type of entity. We included a Suggestions feature to ensure correct word and label alignment, which allows for backtracking and corrections. The agent evaluates the sentence for the presence of the specified entity and assigns appropriate labels in a single-label Beginning/Inside/Outside (BIO) format [40]. Each agent produces a list of named-entity recognition (NER) tags for each sentence, indicating the presence or absence of the entity types.
Summarizing module: After annotating a sentence for each entity type independently, the labels were systematically aggregated into a comprehensive two-dimensional list. Each sub-list contains one or multiple tags for each word in the text, as shown in Figure 2. At this stage, spaces were used as delimiters for checking the alignment of labels, so the labels are word-level, not token-level.
Human-in-the-Loop Review (Annotations)
In total, 3,739 sentences were annotated, each labeled for multi-label token classification. Figure 3(a) depicts the distribution of each type of bias represented in the GUS Dataset, as labeled post-generation by GPT-4o. This was done as part of our quantitative review process, and re-labeling type-of-bias post-generation was done to capture cases where a sentence represents more than one type of bias. While the distribution of biases in the dataset was generally balanced, the distribution of entities was not. Figure 3(b) depicts the distribution of token labels in the annotated GUS dataset. Since each token can be classified with more than one label, the total number of labels is greater than the total number of tokens in the GUS dataset (69,679 tokens). The unbalanced distribution of token labels is inherent to the task, as some entities like UNFAIR are less common than others like O (neutral), and was considered while constructing the training architecture for benchmarking. The GUS dataset is 54.7% statements and 45.3% questions. Having satisfied the human reviewers’ qualitative and quantitative checks, the annotated dataset was deemed ready for downstream tasks, including model training and evaluation.
3.2 Multi-Label Token Classification for Social Bias Detection
Given an input text sequence , where each token represents a word or subword in the sequence, the goal is to classify each token into one or more categories of social bias:
1.
Generalization (GEN)
2.
Unfairness (UNFAIR)
3.
Stereotype (STEREO)
Since tokens can belong to multiple categories simultaneously, this constitutes a multi-label token classification problem. To preserve sequence context and detect nested biases, we employ a B/I/O (Beginning, Inside, Outside) tagging scheme, where:
1.
B-ENTITY: Marks the beginning of a bias-related entity.
2.
I-ENTITY: Marks the continuation of the entity.
3.
O (Outside): Marks tokens that do not belong to any bias category.
We define a multi-hot label matrix for the sequence:
where:
1.
is the number of tokens in the input sequence.
2.
(corresponding to GEN, UNFAIR, and STEREO).
For each token , the model predicts a label vector:
where:
The function maps the input sequence to a predicted label matrix :
To train the model, we minimize the binary cross-entropy (BCE) loss across all tokens and bias categories:
To mitigate class imbalance, we apply focal loss:
The model learns to assign multiple bias labels per token while handling class imbalances effectively using focal loss. The function can be instantiated as a transformer-based model fine-tuned on the GUS dataset.
3.3 Benchmarking Discriminative and Generative Models
In our experiments, we benchmark two distinct classes of models for the multi-label token classification task: discriminative (encoder-only) models and generative (decoder-only) models.
Discriminative Models (Encoder-Only)
For the discriminative approach, we fine-tune pre-trained encoder-only language models. These models extract token embeddings and add a dense layer to map each token to the three target bias categories. The input processing pipeline involves truncating or padding sequences to a fixed length of 128 tokens, aligning word-level labels with subword tokens (with subword tokens inheriting the label of their parent word), and converting entity tags into a matrix. To address class imbalances, we incorporate focal loss during training.
Our evaluation includes transformer-based encoders such as DistilBERT (66M parameters), BERT-base-uncased (110M parameters), and RoBERTa-base (123M parameters), all implemented using the Hugging Face transformers library.
Generative Models (Decoder-Only)
Generative, auto-regressive LLMs are adapted to perform token-level labeling through instruction fine-tuning and few-shot prompting. Due to their sequential generation process, these models are not inherently designed for token-level classification. To accommodate this, we format each input as a chatML prompt that includes a user message (providing instructions, target entities, and the input text) and a system message containing the corresponding true labels from the GUS dataset. This prompt-based format allows us to evaluate the models both quantitatively (by assessing alignment accuracy and standard metrics) and qualitatively (through their reasoning ability).
For parameter-efficient fine-tuning, we employ low-rank adaptations (LoRA) using Unsloth with a LoRA rank of 16 and an of 16; in one instance, a model is quantized to 4-bit precision due to memory constraints. Our experiments with generative models involve Llama 3.3 (70B parameters), Llama 3.2 (3B parameters), Phi 3 Medium (14B parameters), and Phi 3.5 Mini (4B parameters).
Unified Evaluation.
Both discriminative and generative models are benchmarked on the GUS dataset. While encoder-only models offer efficient, direct token-level classification, decoder-only models provide enhanced reasoning capabilities and flexible, prompt-based responses. We evaluate the models using standard token classification metrics (such as precision, recall, and F1 score) along with qualitative assessments of model reasoning.
4 Experimental Settings
4.1 Hardware Setting
All experiments for discriminative (encoder-only) models were conducted on a single NVIDIA T4 GPU with 16GB of memory, while generative auto-regressive (decoder-only) 111In the rest of this document, we use encoder-only to represent discriminative (encoder-only) models and decoder-only to represent generative auto-regressive (decoder-only) LLMs. models were fine-tuned on an A100 with 40GB of memory; both were executed on Google Colab. We also tested decoder-only models for the prompting in few-shots settings with same hardware configurations. The training was implemented using PyTorch and the transformers library, executed on Ubuntu 20.04 with Python 3.8. We used PyTorch Lightning to streamline the training loops and logging mechanisms for encoder fine-tuning, and Unsloth for parameter-efficient fine-tuning of LLMs.
4.2 Evaluation Strategy
Evaluation Data
The dataset developed and used in this study is detailed in Section 3.1. It includes 3,739 annotated samples evenly distributed across types of bias, initially labeled by GPT-4o and subsequently reviewed by human annotators. This dataset was used for fine-tuning both encoder-only and decoder-only models. The data is divided into training, validation, and test sets with an 70-15-15 split ratio by random sample. The BABE [17] dataset was used as an out of distribution corpus in the same settings as we used our dataset.
Evaluation Metrics
This work is a token level classification task so we utilized a variety of metrics that are commonly used for this task in the related works [41] to assess its ability to accurately identify biased entities. These metrics are accuracy-based for evaluating multi-label classification problems.
1.
Precision, Recall, and F1-Score: These metrics were calculated at two levels: individually for each entity class and as a macro-average across all classes. They are defined as follows:
where TP denotes true positives, FP denotes false positives, and FN denotes false negatives.
2.
Hamming Loss: This metric measures the fraction of incorrect labels over all labels in the sequence, accounting for multi-label classification. It is defined as:
where is the total number of tokens multiplied by the number of labels per token, is the true label list for the -th token, is the predicted labels list, and is an indicator function that evaluates whether each true label differs from the prediction labels.
In our evaluation, Hamming loss is effectively similar to accuracy but in the context of multi-label token classification tasks.
Given the imbalanced class distribution in our dataset, we evaluated both the macro-average performance of the model and individual entity-type-level metrics. By treating B- and I- tags as a single entity (e.g., combining B-GEN and I-GEN predictions), we enhance our evaluation of the model’s ability to detect the presence of each biased entity, rather than merely assessing the boundaries. This approach allows us to gain deeper insights into the model’s performance across the diverse classes of bias present in the data, rather than overall accuracy which could show promising metrics even without performing well on our intended entities.
4.3 Evaluation Models and Hyperparameters
Evaluation Strategy
We performed offline evaluation on the test sets using two types of models. The first type includes encoder-only models that were fine-tuned on our specific dataset to adapt to the nuances of our task. The second type comprises decoder-only LLMs, which we explored in two modes: prompting and instruction fine-tuning. We adapted these LLMs to address the challenge of detecting bias at the token-level in textual data. The instruction template for this task would explicitly ask the LLM to identify and classify each token based on its contribution to bias within the text. Below is an example of such an instruction template:
This chat template instructs the LLM to analyze each token in the provided sentence and classify tokens with one or multiple of the GUS entities. The task-specific guidance helps the model to focus on the nuances of language that may convey bias, making it a practical tool for detecting subtle biases in text.
Baseline models Our encoder-only baselines include BERT (trained with focal loss and binary cross-entropy), DistilBERT, and RoBERTa. For decoder-only models, our baselines consist of Llama 3.3, Llama 3.2, Phi 3 Medium, and Phi 3.5 Mini.
We trained our encoder-only multi-label token classification models with seven output classes over 20 epochs. The training process utilized a batch size of 16 and an initial learning rate of . The AdamW optimizer with weight decay was implemented, along with a linear learning rate scheduler featuring a warm-up ratio of 0.1. To handle class imbalance, we used focal loss with and . The classification threshold for all labels was set at 0.5. The original dataset was partitioned into training (75%), validation (15%), and test (10%) splits, ensuring similar distributions of biased entity types across these splits.
For the LLM fine-tuning, a LoRA rank of 16 and of 16 were used, with a dropout of 0. Only Llama 3.3 was quantized to 4-bit, while all other models used 16-bit precision. A temperature of 0.1 was used for the evaluation on a 10% test split of the dataset (the other 90% was used in fine-tuning for 1 epoch). Details of the general hyperparameters and specific QLoRA parameters are provided in Table 3.
| BERT-like Models | LLM Variants |
Models:
BERT-base-uncased, DistilBERT (uncased), RoBERTa (base) |
Models:
Llama3.3-70B, Llama3.2-3B, Phi-3.5-mini-4k, Phi-3-medium-4k |
Learning Rate:
5e-5 with linear scheduler and 10% warm up period |
Learning Rate:
2e-4 with linear scheduler and 5 warm up steps |
Batch Size:
16 |
Batch Size:
8 (Training), 1 (Evaluation) |
Epochs:
20 (early stopping) |
Epochs:
1 |
| Optimizer: AdamW | Optimizer: AdamW |
Weight Decay:
0.01 |
Weight Decay:
0.01 |
Warm-up Steps:
10% of total steps |
Warm-Up Steps: 5 |
Classification Threshold:
0.5 (default for all) |
| LoRA Parameters | |
Parameter: lora_r = 64, lora_alpha = 16, lora_dropout = 0.2
Task Type: CAUSAL_LM Bias: None For Llama 3.3: BitsandBytes: use_4bit = True, bnb_4bit_dtype = float16, bnb_4bit_quant = nf4, use_nested_quant = True |
4.4 Exploratory Data Analysis
We perform an exploratory data analysis on our data and the results are shown in Figure 3. Figures 3(a) and 3(b) depict the distribution of the GUS dataset, in terms of types of bias and entity labels. Figure 3(a) confirms that the dataset is balanced and comprehensive across bias domains. Conversely, Figure 3(b) indicates a class imbalance among entities. Most notably, the neutral entity (O) is present much more often in the dataset than any other token label. Further, some biased entities are much more common than others. For example, stereotypes which are typically multiple tokens long, and are therefore more common in the dataset than unfair tokens, which are usually a single word/descriptor. Entities which are typically longer also contribute to an imbalance between their own B- and I- labels. The domain coverage is balanced, but the inherent label imbalances underline the need for mindfulness during training.
5 Results
In our evaluations, we aim to determine whether encoder-only or decoder-only models are more effective for bias detection in multi-label classification. Our results compare the performance of these model families in recognizing bias for multi-label token classification tasks. The findings are presented and discussed below, with the highest-performing model configuration as GUS-Net: bert-base-uncased, trained with focal loss.
5.1 Comparison between Encoder-only and Decoder-only Models
The analysis presented in the Table 4 focuses on the performance of encoder-only and decoder-only models, specifically fine-tuned on the GUS dataset. This evaluation highlights the effectiveness of each model in bias detection for multi-label classification tasks. Key metrics include Precision, Recall, and F1 score, where higher values indicate better performance. Additionally, the Hamming Loss is reported, with a preference for lower values (below 0.10) to account for class imbalance.
| Model | Hamming Loss | F1 | Precision | Recall | ||
| Encoder-only Models | ||||||
0.05 | 0.80 | 0.82 | 0.77 | ||
| DistilBERT | 0.08 | 0.65 | 0.89 | 0.59 | ||
| RoBERTa | 0.07 | 0.64 | 0.90 | 0.60 | ||
| Nbias (BCE) | 0.06 | 0.68 | 0.93 | 0.63 | ||
| Decoder-only Models | ||||||
| Llama 3.3 | 0.19 | 0.31 | 0.34 | 0.32 | ||
| Llama 3.2 | 0.20 | 0.37 | 0.42 | 0.70 | ||
| Phi 3 Medium | 0.13 | 0.39 | 0.43 | 0.38 | ||
From Table 4, we observe that BERT-base-uncased demonstrates the highest overall performance among encoder-only models, with notable efficiency in balancing precision and recall while maintaining a low Hamming Loss. Conversely, decoder-only models such as Llama 3.2 show higher recall but struggle with overall F1 scores and higher Hamming Loss, indicating potential challenges in handling multi-label classifications. This analysis suggests that encoder-only models may be more suited for this task, especially in contexts where precision in label prediction is critical.
Additionally, Phi 3.5 Mini was evaluated but was unable to generate outputs that could be parsed or aligned with the input. Further observations revealed that while encoder-only models are inherently aligned 100% of the time, the multi-label nature of the task—where each word in the input sequence can have one or many labels—proved challenging for auto-regressive models such as the tested LLMs. For instance, Llama 3.3 output labels were aligned and could be parsed 57.8% of the time, Phi 3 Medium was effective 54.5% of the time, and Llama 3.2 only worked with 14.4% of the inputs, highlighting significant alignment issues with decoder-only models in multi-label token classification tasks.
Key Finding: Encoder-only models, such as BERT model, outperform decoder-only models in terms of both performance metrics and label alignment capabilities, suggesting their superior suitability for complex multi-label classification tasks requiring precise and accurate label predictions.
5.2 Entity-Level Performance of Encoder-only Models
This section evaluates the performance of encoder-only models in identifying biased entities at the token-level within the GUS dataset. We focus on token-level classification to capture occurrences of social bias in individual words and phrases. Table 5 compares encoder-only and decoder-only models.
| Model | Metrics | Macro | Entity-type-based | |||
| Hamming Loss | 0.05 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.80 | 0.74 | 0.61 | 0.90 | 0.95 | |
| Precision | 0.82 | 0.78 | 0.69 | 0.89 | 0.93 | |
| BERT-base (GUS-Net) | Recall | 0.77 | 0.72 | 0.49 | 0.90 | 0.97 |
| Hamming Loss | 0.08 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.65 | 0.66 | 0.14 | 0.86 | 0.92 | |
| Precision | 0.89 | 0.87 | 0.85 | 0.94 | 0.90 | |
| DistilBERT | Recall | 0.59 | 0.53 | 0.08 | 0.80 | 0.94 |
| Hamming Loss | 0.07 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.64 | 0.67 | 0.05 | 0.90 | 0.93 | |
| Precision | 0.90 | 0.85 | 0.89 | 0.94 | 0.92 | |
| RoBERTa | Recall | 0.60 | 0.56 | 0.03 | 0.86 | 0.95 |
| Hamming Loss | 0.06 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.68 | 0.70 | 0.19 | 0.89 | 0.95 | |
| Precision | 0.93 | 0.87 | 0.83 | 0.84 | 0.93 | |
| Nbias (BCE) | Recall | 0.63 | 0.56 | 0.11 | 0.86 | 0.97 |
Result in Table 5 show that BERT-base (GUS-Net) outperforms other encoder models in almost all metrics, particularly excelling in F1 and Recall, with a notably low Hamming Loss, suggesting high accuracy across multiple label types. This model also shows robust performance across different entity types, even those underrepresented in the dataset, such as Unfairness. In contrast, Nbias, while showing high Precision, falls short in Recall, indicating a potential trade-off between capturing all relevant instances of bias and minimizing incorrect bias predictions. DistilBERT and RoBERTa exhibit lower overall performance, particularly in handling the Unfairness entity.
We summarize our observations as: we observe similar performance between GUS-Net and Nbias, indicating a high level of label classification accuracy. GUS-Net shows superior performance, highlighting its effectiveness in detecting the presence of bias, possibly due to focal loss training that focuses on challenging examples. In terns of entity level performance, GUS-Net demonstrates strong performance across different bias entities, particularly in Stereotypes and Generalizations, without compromising on Neutral entity accuracy. In comparison with decoder-only models, the encoder models, like BERT, show significant advantages in multi-label token classification tasks, providing more detailed and effective bias detection compared to decoder-only models, which struggle with label alignment.
Key Finding Encoder models, particularly when incorporating focal loss, excel in complex multi-label token classification, enabling precise and granular bias detection. These models are not only effective in managing inherent class imbalances but also outperform decoder-only models in alignment and resource efficiency during training and inference.
5.3 Entity-Level Performance of Decoder-only Models
We evaluated decoder-only models using two distinct configurations: prompting and instruction-based fine-tuning (IFT). These approaches were designed to enhance the models’ ability to understand and generate appropriate responses based on specific tasks and instructions, particularly in the context of entity recognition tasks within multi-label classification scenarios. Each configuration aimed to leverage the natural strengths of decoder-only architectures in generating coherent and contextually appropriate
Impact of Instruction Fine-tuning
The evaluation of LLMs instruct fine-tuned on the GUS dataset revealed limitations in their ability to effectively handle multi-label token classification tasks. We performed experiment using Llama3.2 and Llama3.3 and Phi 3 Medium to assess their performance on entity-level.
| Model | Metrics | Macro | Entity-type-based | |||
| Llama 3.3 | Hamming Loss | 0.19 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 216/374 | F1 | 0.31 | 0.21 | 0.11 | 0.40 | 0.50 |
| Precision | 0.34 | 0.24 | 0.13 | 0.30 | 0.71 | |
| Recall | 0.32 | 0.19 | 0.10 | 0.62 | 0.39 | |
| Llama 3.2 | Hamming Loss | 0.20 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 54/374 | F1 | 0.37 | 0.23 | 0.04 | 0.57 | 0.63 |
| Precision | 0.42 | 0.22 | 0.10 | 0.78 | 0.57 | |
| Recall | 0.70 | 0.24 | 0.04 | 0.45 | 0.70 | |
| Phi 3 Medium | Hamming Loss | 0.13 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 204/374 | F1 | 0.39 | 0.25 | 0.01 | 0.58 | 0.72 |
| Precision | 0.43 | 0.30 | 0.06 | 0.67 | 0.69 | |
| Recall | 0.38 | 0.22 | 0.01 | 0.51 | 0.76 | |
Results in Table 6 that Llama 3.3 demonstrated the best alignment ability but exhibited lower accuracy metrics compared to Phi 3 Medium. Phi 3 Medium outperformed other models in terms of F1 Score and Hamming Loss, indicating better overall efficiency in label prediction. Both models struggled with the Unfairness category, particularly in terms of F1 Score and Recall, suggesting a difficulty in accurately identifying less frequent labels. Despite higher alignment, the precision and recall for aligned labels varied significantly, reflecting challenges in consistently predicting correct labels across different entity types. The results highlight the inherent challenges faced by decoder-only models in multi-label token classification, especially concerning label alignment and accurate prediction across diverse entity types.
Key Finding The results highlight the inherent challenges faced by decoder-only models in multi-label token classification, especially concerning label alignment and accurate prediction across diverse entity types.
Impact of Few Shot Prompting
For a complementary approach to instruction-based fine-tuning, we evaluated the Llama 3.3 70B model using few-shot prompting. This method aimed to leverage the model’s ability to generalize from limited examples to manage the task of multi-label token classification. However, alignment of labels with input tokens presented similar challenges as observed with fine-tuning. Notably, when provided with no examples of correct labels, the model failed to align any labels, despite clear formatting instructions in the prompt. This underscores the model’s dependency on example driven guidance for accurate performance.
When employing five and ten-shot prompting, the results, although not as comprehensively aligned as those from fine-tuning, showed improved accuracy on the labels that were correctly aligned. In Table 7, we report a Hamming loss of 0.16 with ten-shot prompting, which reflects a 3% improvement over the fine-tuning approach. Despite this improvement, the performance still does not surpass that of encoder-only models, highlighting the inherent limitations of decoder-only models in handling complex multi-label classification tasks effectively without substantial example guidance.
| Examples in Prompt | Metrics | Macro | Entity-type-based | |||
| 5 shot | Hamming Loss | 0.18 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 169/374 | F1 | 0.37 | 0.21 | 0.18 | 0.38 | 0.70 |
| Precision | 0.38 | 0.19 | 0.17 | 0.50 | 0.66 | |
| Recall | 0.37 | 0.24 | 0.20 | 0.31 | 0.75 | |
| 10 shot | Hamming Loss | 0.16 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 148/374 | F1 | 0.38 | 0.23 | 0.19 | 0.36 | 0.75 |
| Precision | 0.40 | 0.23 | 0.18 | 0.49 | 0.69 | |
| Recall | 0.38 | 0.23 | 0.20 | 0.28 | 0.81 | |
Qualitative Analysis
Since the auto-regressive nature of LLMs showed limitations in the multi-label token classification task, we prompted them to identify the entities in example inputs without the necessity to classify specific tokens. From the results in Table 8, we can see many LLMs are effective in identifying subtle or implicit bias, though they appear to have a hard time separating the task of understanding a socially biased context and the task of identifying parts of speech such as generalizations, unfairness, or stereotypes.
| Model | Input Text | Entities Predicted and Reasoning | Observation |
| Llama 3.3 | Young men are always causing trouble. |
Entities: Generalization, unfairness, and stereotype.
Rationale: “Overly broad and inaccurate statement perpetuating negative stereotypes about a specific demographic.” |
Correctly identified the explicit bias against young men, aligning with true labels. |
| Llama 3.2 | The women in the office discussed future improvements for the department. |
Entities: None
Rationale: “The statement highlights women in leadership positions, promoting positive representation in the workplace.” |
Missed the opportunity to recognize the context as non-biased. |
| Phi 3 Medium | He’s really articulate for someone from that neighborhood. |
Entities: Generalization, stereotype.
Rationale: “The statement implies a generalization about people from a certain neighborhood’s ability to be articulate.” |
Correctly identified implicit bias despite the positive sentiment. |
| Phi 3.5 Mini | Black programmers created this software. |
Entities: None.
Rationale: Acknowledges diversity in the tech industry. |
Failed to recognize the unnecessary racial marking as bias. |
Due to the superior performance of encoder-only models in our previous setup, we will now be using GUS-Net with BERT setting in subsequent applications.
5.4 Evaluating Model Performance Against Expert-Annotated Bias
The BABE (Bias Annotations By Experts) dataset [18] is a well-established resource in bias detection, containing a diverse range of biased statements annotated by trained experts. This dataset is valuable as it provides insights into various forms of bias across different demographics and contexts, making it a relevant benchmark for evaluating our model’s performance.
In this analysis, we aimed to compare the normalized number of biased words per sentence in the BABE dataset with the number of positive (non-‘O’) label classifications made by our model (GUS-Net). The normalized number of biased words refers to the count of biased words adjusted for sentence length, allowing for a fair comparison across sentences of varying lengths.
To obtain the normalized number of biased words, we first filtered the training split of the BABE dataset to include only sentences classified as biased. Since our model labels multiple entity types (GEN, UNFAIR, and STEREO) and the BABE dataset does not distinguish between different forms of bias, we adjusted for imbalance by binning the results and using the minimum number of GUS entities found in each bin. The number of biased words from BABE was then normalized by dividing by the sentence length.
The scatter plot in Figure LABEL:fig:baseline_comparison reveals a positive correlation between the normalized number of biased words from the BABE dataset and the normalized minimum number of biased entities predicted by our model. This trend suggests that our model’s understanding of bias aligns well with the definitions established in the BABE dataset, indicating that GUS-Net effectively captures and represents social biases present in the text.
5.5 Ablation Study
We conducted an ablation study on GUS-Net to evaluate the impact of different configurations on the model’s performance. Table 9 presents the macro-average Precision, Recall, F1-score, and Hamming Loss for the following settings: (1) Our proposed GUS-Net model; (2) GUS-Net without GUS dataset: This configuration relies on an existing corpus, BABE [18]. Since there are no token-level annotations for BABE, we used the same annotation pipeline outlined in this paper. (3) GUS-Net without focal loss: In this configuration, we trained the model using the binary cross-entropy (BCE) loss function.
| Metrics | GUS-Net |
| Precision | 0.82 | 0.02 | 0.93 | ||||
| Recall | 0.77 | 0.22 | 0.63 | ||||
| F1-Score | 0.80 | 0.05 | 0.68 | ||||
| Hamming Loss | 0.05 | 0.26 | 0.06 |
From the results in Table 9, we have the following observations:
1.
Our proposed architecture, GUS-Net, outperforms the other configurations across nearly all key performance metrics. Specifically, GUS-Net achieves the highest macro-average Precision (0.82) and F1-Score (0.80), along with the lowest Hamming Loss (0.05), indicating its superior ability to accurately identify and classify entities with minimal misclassifications. The high Precision and F1-Score suggest that GUS-Net effectively reduces false positives while maintaining a strong balance between Precision and Recall.
2.
In contrast, substituting focal loss for BCE resulted in a moderate Precision of 0.65. Upon further inspection of the metrics for each entity individually, we found that the macro-average metrics were distorted by the class imbalance of the ‘O’ tags. Essentially, the model learns to prioritize predicting ‘O’ tags correctly, which detracts from its focus on the new classes of interest. This observation emphasizes the importance of employing a loss function and architecture specifically designed to handle class imbalance, as seen in GUS-Net, ensuring more accurate and reliable model performance.
3.
Interestingly, using the BABE dataset as the underlying corpus for annotation and training yielded poor results. This is likely due to the nature of our test set, which was designed to span various domains, whereas the BABE corpus was gathered specifically from news articles. The domain-specific nature of BABE may limit its effectiveness for generalizing across a broader range of biases.
5.6 Parameter Sensitivity Study
To identify the optimal focal loss parameters, , and , we conducted a sensitivity study by testing various values for each parameter while holding the other constant. As shown in Table 10, we evaluated the performance of the model at different values while keeping fixed at 2. The results indicate that the best-performing value for was , which resulted in improved F1 scores across all entity types. Table 11 shows the influence of parameter while maintaining at . The results reveal that the macro-average F1-Score remained at , indicating that this combination of parameters effectively balances sensitivity and specificity across entity types. Overall, the sensitivity study highlights the importance of tuning the focal loss parameters to improve the model’s performance in identifying various biases. The optimal values used in this paper ( and ) demonstrate the model’s ability to adapt to class imbalances and enhance its performance in detecting biased entities.
| 0.1 | 0.2 | 0.4 | 0.65 | 0.8 | |
| Generalizations F1 | 0.19 | 0.40 | 0.56 | 0.74 | 0.71 |
| Unfairness F1 | 0.01 | 0.14 | 0.35 | 0.61 | 0.54 |
| Stereotypes F1 | 0.60 | 0.81 | 0.83 | 0.90 | 0.83 |
| Neutral F1 | 0.87 | 0.91 | 0.94 | 0.95 | 0.91 |
| Macro Average F1 | 0.42 | 0.57 | 0.67 | 0.80 | 0.75 |
| Hamming Loss | 0.09 | 0.08 | 0.07 | 0.05 | 0.09 |
| 0.5 | 1 | 2 | 3 | 4 | |
| Generalizations F1 | 0.74 | 0.73 | 0.74 | 0.74 | 0.71 |
| Unfairness F1 | 0.55 | 0.48 | 0.61 | 0.57 | 0.57 |
| Stereotypes F1 | 0.90 | 0.89 | 0.90 | 0.88 | 0.87 |
| Neutral F1 | 0.95 | 0.95 | 0.95 | 0.94 | 0.94 |
| Macro Average F1 | 0.78 | 0.76 | 0.80 | 0.78 | 0.77 |
| Hamming Loss | 0.05 | 0.05 | 0.05 | 0.06 | 0.06 |
5.7 Model Validation Example
To demonstrate our model’s labeling capabilities and generalizability, we present a case involving religious bias from the GUS dataset. In Figure LABEL:fig:case(a), we provide an example of a statement that exhibits religious bias, along with the corresponding labels generated by GUS-Net. Figure LABEL:fig:case(b) showcases GUS-Net’s outputs for this case study, illustrating its ability to accurately identify and classify instances of religious bias. The outputs are represented visually, highlighting how the model distinguishes between different types of bias, including Generalizations, Unfairness, and Stereotypes. This example indicates the effectiveness of GUS-Net in generalizing across various forms of bias, reinforcing its potential as a robust tool for bias detection in diverse contexts.
6 Discussion
6.1 Practical and Theoretical Impact
The GUS-Net framework offers practical implications for various domains, including content moderation, social media analysis, and AI-driven auditing tools. By providing a fine-grained, multi-label token classification approach, GUS-Net enables more precise bias identification, reducing the risk of oversimplification seen in traditional binary classification methods. This work is particularly relevant for regulatory and compliance frameworks, where detailed bias categorization is necessary to ensure transparency and accountability.
GUS-Net’s approach to leveraging generative AI for dataset construction demonstrates a scalable solution for addressing data scarcity in bias detection research. This method can be extended to other sensitive NLP applications, such as fairness assessments in hiring platforms or bias mitigation in automated decision-making systems [27]. The findings also reinforce the importance of employing encoder-based models for bias detection tasks, as they offer more reliable token-level classification [42, 43]. This insight informs future model selection strategies for developers and researchers building AI systems for social bias analysis. In future, the integration of explainability techniques, such as attention visualization and interpretability layers, could further enhance the practical applicability of bias detection models. This would allow policymakers and AI practitioners to better understand model predictions and refine mitigation strategies accordingly.
The theoretical impact of the GUS-Net framework lies in its innovative approach to understanding bias detection in NLP. By redefining the problem as a multi-label token classification task, it offers a more nuanced and granular method for identifying biases in text. This approach allows for the detection of multiple, overlapping biases within a single token or phrase, moving beyond traditional binary or single-label classification methods. As a result, GUS-Net provides a deeper and more comprehensive understanding of bias in language, paving the way for more accurate and context-aware bias detection systems in NLP
6.2 Limitations
Just like any study, we also acknowledge a few limitations. First, the use of synthetic data for bias classification raises questions about how well the model can generalize to real-world scenarios. Second, the dataset has an imbalance in entity labels, with neutral tokens (O-labels) being overrepresented. This could skew the model’s learning process, making it more likely to classify tokens as neutral. Although steps were taken to address this, such as using focal loss, other approaches [44] like weighted sampling or creating more balanced datasets could be explored to further improve results.
Third, while the encoder-only models like BERT performed better that decoder only LLMs for multi-label classification problem, the latter could show stronger reasoning abilities [45]. However, auto-regressive models struggled with aligning predicted labels in multi-label tasks [bender2021danger]. It would be useful to combine the strengths of both model families (encoder-only and decoder-only) for the task [46].
Lastly, while the standard NLP benchmarks are used for evaluation, the ethical challenges of bias detection remains subjective [47]. Human biases might still influence the annotation process, and societal changes could shift how bias is defined. In future, it will be a good practice to focus on making training data more fair and ensuring that annotation frameworks are inclusive and adaptable to evolving perspectives.
7 Conclusion
The proposed GUS-Net model addresses limitations in existing bias detection methods by focusing on the nuanced identification of social biases with semantic categories of generalizations, unfairness, and stereotypes. We identify that auto-regressive decoder-only models are poorly suited for multi-label token classification. Conversely, BERT-base-uncased trained on the GUS dataset (GUS-Net) had superior alignment and effectiveness, while allowing for complex label structures that can span multiple words and can be nested/overlapping. GUS-Net approaches bias with three detailed entities, offering a more granular and precise detection of social biases. This enables better insights into the structural components of biased language. Our results demonstrate that GUS-Net performs well at classifying tokens as each of the entities, with a notable strength in detecting stereotypes. In sum, GUS-Net contributes the field of bias detection in NLP by incorporating a fine-grained and multi-faceted view of biased language.
Declaration
Authorship contribution statement
Maximus Powers: Project Lead, Engineering dataset generation, model training & evaluation, Writing drafting, review & editing. Shaina Raza: Advisor, Writing final draft, review & editing. Alex Chang: Writing review & editing, Dataset Quality human-in-the-loop review lead. Umang Mavani: Engineering synthetic corpus pipeline, Dataset Quality human-in-the-loop reviewer. Harshitha Reddy Jonala: Engineering annotation pipeline, Dataset Quality human-in-the-loop reviewer. Ansh Tiwari: Engineering annotation pipeline, Dataset Quality human-in-the-loop reviewer. Hua Wei: Advisor, Writing review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
We thank OpenAI for providing us with API credits under the Researcher Access program.